# A Comprehensive Overview of Fish-Eye Camera Distortion Correction Methods

### v1

@@ 2.1 Equidistant Projection Model
Here, ($\theta$, $\phi$) represents the spherical coordinates on the unit sphere, $r$ is the radial distance from the optical center, ($u$, $v$) represents the normalized image coordinates, and $f$ is the focal length of the fisheye lens.
@@ 2.2 Equiangular Projection Model
Here, ($\theta$, $\phi$) represents the spherical coordinates on the unit sphere, and ($u$, $v$) represents the normalized image coordinates, ranging from 0 to 1.
@@ 3 Distortion Correction Methods
Camera distortion is the alteration of an image’s perspective caused by the camera’s lens, sensor, or other factors. There are several types of distortion, including: Radial distortion, Tangential distortion, as well as Non-linear distortion. The purpose of camera distortion correction is to transform the distorted image captured by the camera into an image that resembles the ideal image produced by a pinhole camera. This correction aims to improve the accuracy of the image, enhance its visual quality, and meet the specific requirements of various applications.
@@ 3.1 Distortion Types
3.1.1 Symmetric Radial Distortions
Symmetric radial distortions are what are typically imagined when discussing image distortion. Often, this type of distortion will be characterized depending on if it is positive (pincushion) or negative (barrel) distortion.
3.1.2 Asymmetric Radial Distortions
Asymmetric radial distortions are radial distortion effects much like the above, but unlike symmetric radial distortion, asymmetric radial distortion characterizes distortion effects that are dependent both on the distance from the image centre as well as how far away the object being imaged is. Asymmetric radial effects are most pronounced in two scenarios:
Cameras with long focal lengths and very-short relative object distances. e.g. a very-near-field telephoto lens that is capturing many objects very close. Observing objects through a medium of high-refraction, or differing refractive indices. e.g. two objects underwater where one is near and one is far away. This type of distortion is typically tricky to visualize, as well as to quantify, because it is dependent on the environment. In most robotic and automated vehicle contexts, asymmetric radial distortion is not a great concern! Why? Well, the difference in distortions depends on the difference in distances between objects. This is usually because of some kind of refractive difference between two objects being imaged, or because the objects are out of focus of the camera (i.e. focal length is too large relative to the object distance).
Neither of the above two scenarios are typical; as such, asymmetric radial distortion is an important aspect of modeling the calibration in applications when these scenarios are encountered.
In most robotic contexts, the primary use for imaging and visual-odometry is done in relatively short ranges with cameras that have short focal lengths, and the primary medium for light to travel through is air. Since there doesn’t tend to be big atmospheric variances between objects that are close, and since light is all traveling through the same medium, there isn’t much of an asymmetric refractive effect to characterize or measure. As a result, this kind of radial distortion isn’t common when calibrating cameras for these kinds of applications. If we can’t measure it, we shouldn’t try to model it!
@@ 3.1.3 Tangential (De-centering) Distortions
Tangential distortion is sometimes also called de-centering distortion, because the primary cause is due to the lens assembly not being centered over and parallel to the image plane. The geometric effect from tangential distortion is not purely along the radial axis. Instead, as can be seen in the figure above, it can perform a rotation and skew of the image plane that depends on the radius from the image centre!
@@ 3.2.1 Brown-Conrady
Brown-Conrady distortion[11] is probably what most think of as the "standard" radial and tangential distortion model. It was first published in 1966 by Brown and Conrady, and has been used in many applications since.
This model characterizes radial distortion as a series of higher order polynomials:
$r=\sqrt{x^{2}+y^{2}}$ (13)
$\delta r=k_{1}r^{3}+k_{2}r^{5}+k_{3}r^{7}+...+k_{n}r^{n+2}$ (14)
In practice, only the $k_{1}$ through $k_{3}$ terms are typically used. For cameras with relatively simple lens assemblies (e.g. only contain one or two lenses in front of the CMOS/CCD sensor), it is often sufficient to just use the $k_{1}$ and $k_{2}$ terms.
To relate this back to our image coordinate system (i.e. $x$ and $y$), we usually need to do some basic trigonometry:
$\delta x_{r}=\sin(\varphi)\delta_{r}=x/r(k_{1}e^{3}+k_{2}r^{5}+k_{3}r^{7})$ (15)
$\delta y_{r}=\cos(\varphi)\delta_{r}=y/r(k_{1}e^{3}+k_{2}r^{5}+k_{3}r^{7})$ (16)
Angential distortion, as characterized by the Brown-Conrady model, is often simplified into the following $x$ and $y$ components. We present these here first as they are probably what most are familiar with:
$\delta x_{t}=2p_{2}xy+p_{1}(r^{2}+2x^{2})$ (17)
$\delta y_{r}=p_{2}(r^{2}+2y^{2})+2p_{1}xy$ (18)
This actually derives from an even-power series much like the radial distortion is an odd-power series. The full formulation is a solution to the following:
$\delta t=P(r)\cos(\varphi-\varphi_{0})$ (19)
Where p(r) is our de-centering distortion profile function, is the polar angle of the image plane coordinate, and is the angle to the axis of maximum tangential distortion (i.e. zero radial distortion). Expanding this into the general parameter set we use today is quite involved (read the original Brown paper!), however this will typically take the form:
$\delta x_{t}=[p_{1}(r^{2}+2x^{2})+2p_{2}xy](1+p_{3}r^{2}+p_{4}r^{4}+p_{5}r^{6}+...)$ (20)
$\delta y_{r}=[2p_{1}xy+p_{2}(r^{2}+2y^{2})](1+p_{3}r^{2}+p_{4}r^{4}+p_{5}r^{6}+...)$ (21)
Because tangential distortion is usually small, we tend to approximate it using only the first two terms. It is rare for de-centering to be so extreme that our tangential distortion requires higher order terms because that would mean that our lens is greatly de-centered relative to our image plane. In most cases, one might ask if their lens should simply be re-attached in a more appropriate manner.
@@ 3.2.2 Kannala-Brandt
Almost a century later (2006, from the original Conrady paper in 1919), Juho Kannala and Sami Brandt[12] published their own paper on lens distortions. The main contribution of this paper adapts lens distortion modeling to be optimized for wide-angle, ultra wide-angle, and fish-eye lenses. Brown Conrady’s modeling was largely founded on the physics of Seidel aberrations, which were first formulated around 1867 for standard lens physics of the time, which did not include ultra wide and fish-eye lenses.
The primary difference that most folks will notice using this model lies in symmetric radial distortion. Rather than characterizing radial distortion in terms of how far a point is from the image centre (the radius), Kannala-Brandt characterizes distortion as a function of the incidence angle of the light passing through the lens. This is done because the distortion function is smoother when parameterized with respect to this angle ( ), which makes it easier to model as a power-series:
$\theta=\arctan(r,f)$ (22)
$\delta r=k_{1}\theta^{2}+k_{2}\theta^{4}+k_{3}\theta^{6}+...$ (23)
Above, we’ve shown the formula for $\theta$ when using perspective projection, but the main advantage of the Kannala-Brandt model is that it can support different kinds of projection by swapping our formula for $\theta$, which is what makes the distortion function smoother for wide-angle lenses. Kannala-Brandt also aims to characterize other radial (such as asymmetric) and tangential distortions. This is done with the following additional parameter sets:
$\delta r_{o}ther=(l_{1}\theta+l_{2}\theta^{3}+l_{3}\theta^{5})(i_{1}\cos\phi+i_{2}\cos(2\phi)+i_{4}\sin(2\phi)+...)$ (24)
@@
Overall, this results in a 23 parameter model! This is admittedly overkill, and the original paper claims as much. These models, unlike the symmetric radial distortion, are an empirical model derived by fitting an N-term Fourier series to the data being calibrated. This is one way of characterizing it, but over-parameterizing our final model can lead to poor repeatability of our final estimated parameters. In practice, most systems will characterize Kannala-Brandt distortions purely in terms of the symmetric radial distortion, as that distortion is significantly larger in magnitude and will be the leading kind of distortion in wider-angle lenses.
@@ 3.4 Direct Methods
Direct methods for fish-eye camera distortion correction involve the detection and analysis of specific features or patterns in the image to estimate the distortion parameters. These methods typically rely on the relationships between the distorted image coordinates and the undistorted object coordinates. Techniques such as RANSAC (Random Sample Consensus) can be used to robustly estimate the distortion parameters from the feature correspondences. Direct methods are advantageous in scenarios where calibration data or a priori knowledge about the distortion model is not available.
@@ 3.4.2 Latitude-Longitude Mapping Method
The Latitude-Longitude Mapping Method is a technique used for transforming a fisheye image into a panoramic or equirectangular projection. This method involves mapping the distorted fisheye pixels to their corresponding latitude and longitude coordinates on a spherical or cylindrical coordinate system, and then projecting them onto a 2D equirectangular grid.
@@ 3.4.3 Panorama Mapping Method
The Panorama Mapping Method is a technique used for fisheye image rectification and distortion correction. It aims to transform a distorted fisheye image into a rectilinear image, which has straight lines and a more natural perspective. The method involves mapping the distorted fisheye pixels to their corresponding locations in the rectilinear image.
@@ 3.5.1 Converlutional Meural Metworks(CNNs)
Convolutional Neural Networks (CNNs) have been extensively employed for fisheye camera distortion correction. These networks consist of multiple convolutional layers that extract hierarchical features from the input images. By training CNNs on a large dataset of distorted and undistorted fisheye image pairs, they can learn the underlying patterns and relationships to predict the distortion-free version of a given fisheye image. (Rong et. al., 2016)[19] intends to employ CNNs (Convolutional Neural Networks), to achieve radial distortion correction. Inspired by the growing availability of image dataset with non-radial distortion (Rong et. al., 2016) propose a framework to address the issue by synthesizing images with radial distortion for CNNs. To this end (Lu et. al., 2017)[20] propose a fast level set model-based method for intensity inhomogeneity correction and a spectral properties-based color correction method to overcome these obstacles. In contrast to conventional approaches, the proposed model integrates a new signed energy force function that can detect contours at weak or blurred edges efficiently. (Borkar et. al., 2019)[21] evaluate the effect of image distortions like Gaussian blur and additive noise on the activations of pre-trained convolutional filters. (Borkar et. al., 2019) propose a metric to identify the most noise susceptible convolutional filters and rank them in order of the highest gain in classification accuracy upon correction. Radial lens distortion often exists in images taken by commercial cameras, which does not satisfy the assumption of pinhole camera model. They generated images with a large number of images of high variation of radial distortion, which can be well exploited by deep CNN with a high learning capacity, and reach the state-of-the-art results. (Shi et. al., 2018)[22] claim that a weight layer with inverted foveal models can be added to these existing CNNs methods for radial distortion correction. Convolutional neural networks (CNNs) have been widely used for road scene understanding in the last few years with great success. (Arsenali et. al., 2019)[23] propose RotInvMTL: a multi-task network (MTL) to perform joint semantic segmentation, boundary prediction, and object detection directly on raw fisheye images. An attempt to optimize a CNN-based detector for fisheye cameras was made, taking into consideration the barrel distortion, which complicates the object detection (Goodarzi et. al., 2019)[24]. The obtained result proves that fisheye augmentation can considerably advance a CNN-based detector’s performance on fisheye images in spite of the distortion. (Sáez et. al., 2019)[25] present a methodology that provides real-time semantic segmentation on fisheye cameras leveraging only synthetic images. (Sáez et. al., 2019) propose some Convolutional Neural Networks(CNN) architectures based on Efficient Residual Factorized Network(ERFNet) that demonstrate notable skills handling distortion and a new training strategy that improves the segmentation on the image borders. (Vasiljevic et. al., 2020)[26] show that self-supervision can be used to learn accurate depth and ego-motion estimation without prior knowledge of the camera model. Inspired by the geometric model of Grossberg and Nayar, (Vasiljevic et. al., 2020) introduce Neural Ray Surfaces (NRS), convolutional networks that represent pixel-wise projection rays, approximating a wide range of cameras. The strong radial distortion breaks the translation invariance inductive bias of Convolutional Neural Networks. (Ramachandran et. al., 2022)[27] provide a detailed analysis on the competition which attracted the participation of 120 global teams and a total of 1492 submissions. The fisheye image has a severe geometric distortion which may interfere with the stage of image registration and stitching. In the stage of fisheye image correction (Hao et. al., 2023)[28] propose the Attention-based Nonlinear Activation Free Network (ANAFNet) to deblur fisheye images corrected by Zhang calibration method.
@@ 3.5.2 Generative Adversarial Networks(GANs)
(Li et. al., 2012)[29] present a novel embedded real-time fisheye image distortion correction algorithm with application in IP network camera. A fast and simple distortion correction method is introduced based on Midpoint Circle Algorithm (MCA) which aims to determine the pixel positions along a circle circumference based on incremental calculation of decision parameters. Each lens is calibrated separately and interior/relative orientation parameters (IOPs and ROPs) of the camera are determined on the basis of designed calibration network on the central and side images captured by the aforementioned lenses (Aghayari et. al., 2017)[30]. Designed calibration network is considered as a free distortion grid and applied to the measured control points in the image space as correction terms by means of bilinear interpolation. Deep learning techniques have become popular for performing camera model identification. To expose weaknesses in these methods (Chen et. al., 2018)[31] propose a new anti-forensic framework that utilizes a generative adversarial network (GAN) to falsify an image’s source camera model. (Nikonorov et. al., 2019)[32] present a new end-to-end framework applying two convolutional neural networks (CNNs) to reconstruct images captured with multilevel diffractive lenses (MDLs). The generative adversarial network (GAN) is first used to remove image-wise color distortion, while a patch-wise network is then used to apply chromatic deblur. Correction of the distortion of images is crucial in many computer vision applications. (Liao et. al., 2020)[33] present distortion rectification generative adversarial network (DR-GAN), a conditional generative adversarial network (GAN) for automatic radial DR. To the best of the knowledge, this is the first end-to-end trainable adversarial framework for radial distortion rectification. (Gallego et. al., 2020)[34] present a network architecture with parallel convolutional neural networks (CNN) for removing perspective distortion in images. While other works generate corrected images through the use of generative adversarial networks or encoder-decoder networks, (Gallego et. al., 2020) propose a method wherein three CNNs are trained in parallel, to predict a certain element pair in the 3*3 transformation matrix. Generative adversarial networks (GANs) have been implemented to convert IR images into RGB images for enriching semantic information. (Zhang et. al., 2021)[35] study wggan: a wavelet-guided generative adversarial network for thermal image translation. A wavelet-guided generative adversarial network (WGGAN) is proposed to address the problem. In order to improve the quality of low-light image (Zhang et. al., 2021) propose a Heterogenous low-light image enhancement method based on DenseNet generative adversarial network. Secondly, the feature map from low light image to normal light image is learned by using the generative adversarial network. (Thapa et. al., 2021)[36] present the distortion-guided network (DG-Net) for restoring distortion-free underwater images. (Thapa et. al., 2021) then use a generative adversarial network guided by the distortion map to restore the sharp distortion-free image. (Luo et. al., 2021)[37] propose an unsupervised deep convolutional network that takes rectified stereo image pairs as input and outputs corresponding dense disparity maps. Second, the left and right images, which are reconstructed based on the input image pair and corresponding disparities as well as the vertical correction maps, are regarded as the outputs of the generative term of the generative adversarial network (GAN).
3.5.3 Encoder-Decoder Architectures
Encoder-decoder architectures, such as U-Net and its variants, have proven effective for fisheye camera distortion correction. These architectures consist of an encoder network that captures the high-level features of the distorted fisheye image and a decoder network that reconstructs the undistorted image from the encoded features. The encoder-decoder structure allows for the preservation of spatial information during the distortion correction process.

### v2

@@ 2.1 Equidistant Projection Model
Here, ($\theta$, $\varphi$) represents the spherical coordinates on the unit sphere, $r$ is the radial distance from the optical center, ($u$, $v$) represents the normalized image coordinates, and $f$ is the focal length of the fisheye lens.
@@ 2.2 Equiangular Projection Model
Here, ($\theta$, $\varphi$) represents the spherical coordinates on the unit sphere, and ($u$, $v$) represents the normalized image coordinates, ranging from 0 to 1.
@@ 3 Distortion Correction Methods
Camera distortion is the alteration of an image’s perspective caused by the camera’s lens, sensor, or other factors. There are several types of distortion, including: Radial distortion, Tangential distortion, as well as Non-linear distortion.
The purpose of camera distortion correction is to transform the distorted image captured by the camera into an image that resembles the ideal image produced by a pinhole camera. This correction aims to improve the accuracy of the image, enhance its visual quality, and meet the specific requirements of various applications.
@@ 3.1 Distortion Types
3.1.1 Radial Distortions
Barrel Distortion: In barrel distortion, straight lines near the edges of the image appear to curve outward, giving the image a barrel-like or fisheye effect. This distortion occurs because light rays passing through the outer parts of the lens bend more than those passing through the center.
Pincushion Distortion: Pincushion distortion is the opposite of barrel distortion. In pincushion distortion, straight lines near the edges of the image appear to curve inward, resembling the shape of a pincushion. This occurs when light rays passing through the center of the lens bend more than those passing through the outer parts.
@@ 3.1.2 Tangential (De-centering) Distortions
Tangential distortion, also known as de-centering distortion, arises when the lens assembly is not perfectly aligned with the image plane. This misalignment causes the image to be distorted in a non-radial manner, resulting in effects such as rotation and skewing. Unlike radial distortion, tangential distortion is not purely along the radial axis but varies with the radius from the image center.
@@ 3.2.1 Brown-Conrady
Brown-Conrady distortion[11] is a widely used lens distortion model in computer vision and photogrammetry. The model is primarily employed to correct the geometric distortions caused by camera lenses, which can occur due to various factors such as imperfections in lens manufacturing, non-linearities in the imaging process, and physical limitations of the lens design.
The key components of the Brown-Conrady model include two aspects.
Radial distortion (pincushion and barrel distortion): Light rays bend more at points farther from the center of the lens than those closer to the center. This is expressed mathematically as:
$\delta r_{s}=k_{1s}r^{3}+k_{2s}r^{5}+k_{3s}r^{7}+...+k_{ns}r^{n+2}$ (13)
where $r$ is the distance from the center of the image, and $k_{1s}$, $k_{2s}$, $k_{3s}$, etc., are the radial distortion coefficients.
Tangential distortion: The lens is not perfectly aligned with the image plane, meaning there is an angle error between the sensor assembly and the lens. This is expressed mathematically as:
$\delta x=2p_{1}xy+p_{2}(r^{2}+2x^{2})$ (14)
$\delta y=p_{1}(r^{2}+2y^{2})+2p_{2}xy$ (15)
where $x$ and $y$ are the image coordinates, and $p_{1}$ and $p_{2}$ are the tangential distortion coefficients. After merging the mathematical models of both types of distortion, they can be represented by the following formula:
$x_{corrected}=x_{distorted}(1+k_{1}r^{2}+k_{2}r^{4}+k_{3}r^{6}+\dots)+(2p_{1}x_{distorted}y_{distorted}+p_{2}(r^{2}+2x_{distorted}^{2}))$ (16)
$y_{corrected}=y_{distorted}(1+k_{1}r^{2}+k_{2}r^{4}+k_{3}r^{6}+\dots)+(p_{1}(r^{2}+2y_{distorted}^{2})+2p_{2}x_{distorted}y_{distorted})$ (17)
where $x_{distorted}$ and $y_{distorted}$ are the distorted image coordinates, and $x_{corrected}$ and $y_{corrected}$ are the corrected image coordinates. The parameters $k_{1}$, $k_{2}$, $k_{3}$, $p_{1}$, and $p_{2}$ are the distortion coefficients that need to be estimated during the calibration process.
@@ 3.2.2 Kannala-Brandt
The Kannala-Brandt distortion model[12] is also a mathematical model used to describe radial lens distortion. It’s an extension of the classical Brown-Conrady distortion model and provides a more accurate representation of distortion, particularly for wide-angle lenses and fisheye lenses.
Kannala-Brandt summarized four models for fisheye imaging, along with the pinhole imaging model, and compared the distortion curves of each model.
The perspective projection of a pinhole camera can be described by the following formula:
$x=ftan\theta$ (18)
The four fisheye projection models are:
  • Stereographic projection
    $x=2f\tan\frac{\theta}{2}$ (19)
  • Equidistant projection
    $x=f\theta$ (20)
  • Equiangular projection
    $x=2f\sin\frac{\theta}{2}$ (21)
  • Orthographic projection
    $x=f\sin\theta$ (22)
Figure 1: (a) Projections with f=1 ((i) to (v) are representing the five projection models mentioned above). (b) Fish-eye camera model.[12]
$x=r(\theta)\cos\varphi$ (23)
$y=r(\theta)\sin\varphi$ (24)
The Kannala-Brandt model represents radial distortion using a polynomial function. Unlike the Brown-Conrady model, which typically uses a fixed number of polynomial terms (e.g., $k_{1}$, $k_{2}$, $k_{3}$), the Kannala-Brandt model employs a more flexible approach. It uses a radial polynomial with terms up to a certain order (typically denoted by $n$), allowing for a more accurate representation of complex distortion patterns. To obtain a widely applicable, flexible model, they propose to use two distortion terms as follows: One distortion term acts in the radial direction.
$\delta r_{s}=(l_{1}\theta+l_{2}\theta^{3}+l_{3}\theta^{5})(i_{1}\cos\phi+i_{2}\cos(2\phi)+i_{4}\sin(2\phi)+...)$ (25)
The other distortion term acts in the tangential direction.
@@
The complete camera model comprises 23 parameters, denoted as $p_{23}$ hereafter. Given the high flexibility of the asymmetric component of the model, there are situations where utilizing a simplified camera model may be preferable to prevent overfitting.
@@ 3.3.1 Corner Detection and Rectification
(Sixian Chan et al., 2016)[15] present an enhanced automatic detection method for checkerboards, aiming to circumvent the constraints and user intervention typically associated with conventional methods. They evaluate a state-of-the-art corner detection technique, analyzing its advantages and limitations. To address the identified shortcomings, they develop an adaptive automatic corner detection algorithm.
(Zhang et al., 2020)[16] proposes an optimized calibration method tailored for ultra-wide dual-band cameras, leveraging a specifically designed thermal radiation checkerboard. The optimizations focus on subpixel location, center location, and mapping relationship refinement. Subpixel corner detection involves utilizing a variable quadrilateral template on edge gradients, with the minimum orthogonal value calculated to precisely locate the corners. Center location is achieved through Gaussian fitting, iteratively selecting the bottom point as the potential center.
@@ 3.3.3 Line Detection and Rectification
A groundbreaking study[20] introduced the concept that straight line segments in the physical world should retain their linearity even after being projected by a fisheye lens. Building upon this principle, Bukhari et al.[21] proposed the utilization of an extended Hough transform of lines to correct radial distortions.
(M. Zhang et. al., 2015)[22] presents a novel algorithm for rectifying fisheye images in the undistorted perspective image plane by incorporating line constraints. They introduced a novel Multi-Label Energy Optimization (MLEO) method for merging short circular arcs with similar circular parameters and selecting long circular arcs to facilitate camera rectification.
(Zhu-Cun Xueet. al., 2020)[23] introduced a pioneering approach termed Line-aware Rectification Network (LaRecNet), aimed at tackling the challenge of fisheye distortion rectification. LaRecNet leverages the fundamental principle that straight lines in 3D space should retain their linearity in image planes.
@@ 3.3.4 Optical Flow-based Methods
Existing methods often overlook the temporal correlation between frames, leading to temporal jitter in the corrected video. To tackle this issue, researchers propose a temporal weighting scheme that gradually reduces the weight of frames, resulting in a more coherent global optical flow (Shangrong Yang, et al., 2023)[25]. By leveraging inter-frame optical flow, researchers can better perceive local spatial deformations in fisheye videos. Additionally, spatial deformation is derived using flows from both fisheye and distortion-free videos, enhancing the local accuracy of predictions. However, correcting each frame independently disrupts temporal correlation. To address this, a temporal deformation aggregator is introduced to reconstruct deformation correlation between frames and provide a reliable global feature[25].
@@ 3.4 Direct Methods
Direct methods for fish-eye camera distortion correction involve the detection and analysis of specific features or patterns in the image to estimate the distortion parameters. These methods typically rely on the relationships between the distorted image coordinates and the undistorted object coordinates. Techniques such as RANSAC (Random Sample Consensus)[26] can be used to robustly estimate the distortion parameters from the feature correspondences. Direct methods are advantageous in scenarios where calibration data or a priori knowledge about the distortion model is not available.
@@ 3.4.2 Latitude-Longitude Mapping Method
The Latitude-Longitude Mapping Method was was first proposed by Kum et al. [27]. This process resembles the mapping of the Earth’s surface onto a plane. Initially, each image point is mapped onto a sphere within the latitude and longitude coordinate framework. Subsequently, the points along each meridian share the same abscissa value based on the latitude and longitude distribution, with greater distortion occurring as longitude values increase. Similarly, points along each latitude share the same ordinate value. However, the upper and lower regions of the fish-eye image often exhibit significant stretching and distortion due to excessive correction applied to the original image.
@@ 3.4.3 Panorama Mapping Method
The Panorama Mapping Method is primarily employed to achieve a complete 360° panoramic vista[28]. It entails expanding the field of view to capture the surrounding environment from all angles..
@@ 3.5.1 Converlutional Meural Metworks(CNNs)
Convolutional Neural Networks (CNNs) have been extensively employed for fisheye camera distortion correction. These networks consist of multiple convolutional layers that extract hierarchical features from the input images. By training CNNs on a large dataset of distorted and undistorted fisheye image pairs, they can learn the underlying patterns and relationships to predict the distortion-free version of a given fisheye image.
(Rong et. al., 2016)[29] intends to employ CNNs (Convolutional Neural Networks), to achieve radial distortion correction. Inspired by the growing availability of image dataset with non-radial distortion (Rong et. al., 2016) propose a framework to address the issue by synthesizing images with radial distortion for CNNs.
(Borkar et. al., 2019)[30] evaluate the effect of image distortions like Gaussian blur and additive noise on the activations of pre-trained convolutional filters. (Borkar et. al., 2019) propose a metric to identify the most noise susceptible convolutional filters and rank them in order of the highest gain in classification accuracy upon correction. Radial lens distortion often exists in images taken by commercial cameras, which does not satisfy the assumption of pinhole camera model. They generated images with a large number of images of high variation of radial distortion, which can be well exploited by deep CNN with a high learning capacity.
(Shi et. al., 2018)[31] claim that a weight layer with inverted foveal models can be added to these existing CNNs methods for radial distortion correction.
(Arsenali et. al., 2019)[32] propose RotInvMTL: a multi-task network (MTL) to perform joint semantic segmentation, boundary prediction, and object detection directly on raw fisheye images. An attempt to optimize a CNN-based detector for fisheye cameras was made, taking into consideration the barrel distortion, which complicates the object detection (Goodarzi et. al., 2019)[33]. The obtained result proves that fisheye augmentation can considerably advance a CNN-based detector’s performance on fisheye images in spite of the distortion. (Vasiljevic et. al., 2020)[34] show that self-supervision can be used to learn accurate depth and ego-motion estimation without prior knowledge of the camera model.
(Vasiljevic et. al., 2020)[34] introduce Neural Ray Surfaces (NRS), convolutional networks that represent pixel-wise projection rays, approximating a wide range of cameras. (Ramachandran et. al., 2022)[35] provide a detailed analysis on the competition which attracted the participation of 120 global teams and a total of 1492 submissions. The fisheye image has a severe geometric distortion which may interfere with the stage of image registration and stitching. In the stage of fisheye image correction (Hao et. al., 2023)[36] propose the Attention-based Nonlinear Activation Free Network (ANAFNet) to deblur fisheye images corrected by Zhang calibration method.
(Yang et. al., 2021)[37] propose Progressively Complementary Network which focuses on developing an interpretable correction mechanism for fisheye distortion rectification networks and introduces a feature-level correction scheme. It incorporates a correction layer within skip connections and utilizes appearance flows from various layers to pre-correct image features. As a result, the decoder can more effectively reconstruct the final output using the remaining distortion-free information.
@@ 3.5.2 Generative Adversarial Networks(GANs)
(Li et. al., 2012)[38] present a novel embedded real-time fisheye image distortion correction algorithm with application in IP network camera. A fast and simple distortion correction method is introduced based on Midpoint Circle Algorithm (MCA) which aims to determine the pixel positions along a circle circumference based on incremental calculation of decision parameters.
In (Aghayari et. al., 2017)[39], individual lens calibration is conducted, and the interior/relative orientation parameters (IOPs and ROPs) of the camera are established using a tailored calibration network applied to both central and side images captured by these lenses. Designed calibration network is considered as a free distortion grid and applied to the measured control points in the image space as correction terms by means of bilinear interpolation.
(Nikonorov et. al., 2019)[40] present a new end-to-end framework applying two convolutional neural networks (CNNs) to reconstruct images captured with multilevel diffractive lenses (MDLs).
(Liao et. al., 2020)[41] present distortion rectification generative adversarial network (DR-GAN), a conditional generative adversarial network (GAN) for automatic radial DR. To the best of the knowledge, this is the first end-to-end trainable adversarial framework for radial distortion rectification. ß (Gallego et. al., 2020)[42] propose a method wherein three CNNs are trained in parallel, to predict a certain element pair in the 3*3 transformation matrix. In order to improve the quality of low-light image (Zhang et. al., 2021)[43] propose a Heterogenous low-light image enhancement method based on DenseNet generative adversarial network. Secondly, the feature map from low light image to normal light image is learned by using the generative adversarial network.
(Thapa et. al., 2021)[44] present the distortion-guided network (DG-Net) for restoring distortion-free underwater images. (Thapa et. al., 2021)[44] then use a generative adversarial network guided by the distortion map to restore the sharp distortion-free image. (Luo et. al., 2021)[45] propose an unsupervised deep convolutional network that takes rectified stereo image pairs as input and outputs corresponding dense disparity maps. Second, the left and right images, which are reconstructed based on the input image pair and corresponding disparities as well as the vertical correction maps, are regarded as the outputs of the generative term of the generative adversarial network (GAN).
3.5.3 Vision Transformer(ViT)
Inspired by the significant success of transformer architectures in natural language processing (NLP), researchers have recently extended their application to computer vision (CV) tasks. Vision Transformers (ViTs) have been proposed for fisheye camera distortion correction, leveraging the self-attention mechanism to capture long-range dependencies in the fisheye images. By training ViTs on a large dataset of distorted and undistorted fisheye images, they can learn the complex relationships between the distorted and undistorted pixels and perform distortion correction effectively.
Constrained by a fixed receptive field, existing methods have not fully exploited the global distribution and local symmetry of distortion. To address this, Fishformer[46] was introduced, processing fisheye images as sequences to enhance both global and local perception. The Transformer was fine-tuned based on the structural properties of fisheye images.
In fisheye images, distinct distortion patterns are regularly distributed across the image plane, regardless of the visual content, offering valuable cues for rectification. To leverage these cues effectively, SimFIR[47] is introduced as a straightforward framework for fisheye image rectification through self-supervised representation learning. The technical process involves splitting a fisheye image into multiple patches and extracting their representations using a Vision Transformer (ViT).
4 Experiments
4.1 Datasets
Currently, there is no universally accepted fisheye distortion correction dataset. Therefore, the mainstream approach at present is to utilize generated datasets.
(Yang et. al., 2021)[46] select two completely different types of datasets:(1)Place2 dataset[48] and (2)CelebA dataset[49]. They crop the image to be circular, and then leverage the four-parameters polynomial model to transform the normal image, generating fisheye images.
(Fan et. al., 2021)[50] evaluate the proposed SIR-Net on the synthetic datasets which contains 4 different datasets:ADE20k[51], WireFrame[52], COCO2017[53], and Place365 dataset[48]. They used three distortion models,i.e., FOV[54], Division Model[55], EquiDistant distortion model[56], and synthesized the dataset for each distortion model separately.
(Nobuhiko Wakai et. al., 2021)[57] used two large-scale datasets of outdoor panoramas called the StreetLearn dataset (Manhattan 2019 subset)[58] and the SP360 dataset[59]. Figure 2 illustrates the random distribution of the training set generated by capturing image patches using camera models with the maximum incident angle $\eta_{max}$.
Figure 2: Distribution of the camera parameters for dataset.[57]
4.2 Evaluation Metrics
Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM)[60] stand as the two predominant evaluation metrics in image processing. PSNR provides an effective measure of the image’s detailed quality, while SSIM offers an intuitive assessment of the image’s structural integrity.